OpenStudy (anonymous):

Find the center, vertices, and foci of the ellipse with equation x^2/100 + y^2/64= 1.

OpenStudy (anonymous):

haven't done ellipises yet

OpenStudy (jdoe0001):

mooor, do you have your ellipse formula? does it resemble what you have?

OpenStudy (anonymous):

no :/

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

ah I'm still a bit confused :/

OpenStudy (luigi0210):

What jdoe pretty much showed in the image Confused about what?

OpenStudy (jdoe0001):

well, check 100 and 64, the denominators, can you get a square root off them?

OpenStudy (jdoe0001):

\[\frac{x^2}{100}+\frac{y^2}{64}=1 \implies\\ \frac{(x^2-0)}{10^2}+\frac{(y^2-0)}{8^2}=1 \\ center\ of\ the\ ellipse\ is\ at\ (h,k)\ from\ the\ formula \\ \\ foci = \sqrt={a^2+b^2} \implies \sqrt{10^2+8^2}\]

OpenStudy (jdoe0001):

\[foci = \sqrt{a^2+b^2} \implies \sqrt{10^2+8^2}\]

OpenStudy (jdoe0001):

\[\frac{x^2}{100}+\frac{y^2}{64}=1 \implies\\ \frac{(x-0)^2}{10^2}+\frac{(y-0)^2}{8^2}=1 \\ center\ of\ the\ ellipse\ is\ at\ (h,k)\ from\ the\ formula \\ \\ foci = \sqrt{a^2+b^2} \implies \sqrt{10^2+8^2}\] there, better :)

OpenStudy (jdoe0001):

\[distance\ from\ center = \sqrt{a^2+b^2} \implies \sqrt{10^2+8^2}\\ foci = (h\ or\ k) + \sqrt{10^2+8^2}\]

OpenStudy (jdoe0001):

once you find your center (h,k), move THAT MUCH to find your foci, moving always over the major axis, of course the major axis will be the one with the bigger number(denominator)

OpenStudy (anonymous):

so from theese 4 choices: Center: (0, 0); Vertices: (-10, 0), (10, 0); Foci: (-8, 0), (8, 0) Center: (0, 0); Vertices: (0, -10), (0, 10); Foci: (0, -8), (0, 8) Center: (0, 0); Vertices: (-10, 0), (10, 0); Foci: (-6, 0), (6, 0) Center: (0, 0); Vertices: (0, -10), (0, 10); Foci: (0, -6), (0, 6) it would be the first, correct? :)

OpenStudy (anonymous):

so from theese 4 choices: Center: (0, 0); Vertices: (-10, 0), (10, 0); Foci: (-8, 0), (8, 0) Center: (0, 0); Vertices: (0, -10), (0, 10); Foci: (0, -8), (0, 8) Center: (0, 0); Vertices: (-10, 0), (10, 0); Foci: (-6, 0), (6, 0) Center: (0, 0); Vertices: (0, -10), (0, 10); Foci: (0, -6), (0, 6) it would be the first, correct? :)

OpenStudy (anonymous):

so from theese 4 choices: Center: (0, 0); Vertices: (-10, 0), (10, 0); Foci: (-8, 0), (8, 0) Center: (0, 0); Vertices: (0, -10), (0, 10); Foci: (0, -8), (0, 8) Center: (0, 0); Vertices: (-10, 0), (10, 0); Foci: (-6, 0), (6, 0) Center: (0, 0); Vertices: (0, -10), (0, 10); Foci: (0, -6), (0, 6) it would be the first, correct? :)

OpenStudy (anonymous):

so from theese 4 choices: Center: (0, 0); Vertices: (-10, 0), (10, 0); Foci: (-8, 0), (8, 0) Center: (0, 0); Vertices: (0, -10), (0, 10); Foci: (0, -8), (0, 8) Center: (0, 0); Vertices: (-10, 0), (10, 0); Foci: (-6, 0), (6, 0) Center: (0, 0); Vertices: (0, -10), (0, 10); Foci: (0, -6), (0, 6) it would be the first, correct? :)

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@009infinity

OpenStudy (anonymous):

@009infinity

OpenStudy (e.mccormick):

OK, so you have the work and just need to onfirm at this point, right?

OpenStudy (e.mccormick):

I'm looking for a better description of the foci formula for you. I think that is wrong in that answer, but want to be sure.

OpenStudy (e.mccormick):

When you find the center, which is just \((0,0)\) in this case, you need to find out how far to go out from the center to be at the foci. This takes the form of: (Radius of Major)\(^2\) - (Radius of Minor)\(^2\) = (Focal Distance)\(^2\) This is most commony shortened to: \(a^2-b^2=c^2\) Here is a reference to that: http://www.dummies.com/how-to/content/how-to-graph-an-ellipse.html So I am not sure where that one person got: \(\sqrt{10^2+8^2}\), that would put the foci at \(\pm\sqrt{164}\), whicht hey are not.